# Neural MJD: Neural Non-Stationary Merton Jump Diffusion for Time Series Prediction

### v1

@@ Abstract
While deep learning methods have achieved strong performance in time series prediction, their black-box nature and inability to explicitly model underlying stochastic processes often limit their generalization to non-stationary data, especially in the presence of abrupt changes. In this work, we introduce Neural MJD, a neural network based non-stationary Merton jump diffusion (MJD) model. Our model explicitly formulates forecasting as a stochastic differential equation (SDE) simulation problem, combining a time-inhomogeneous Itô diffusion to capture non-stationary stochastic dynamics with a time-inhomogeneous compound Poisson process to model abrupt jumps. To enable tractable learning, we introduce a likelihood truncation mechanism that caps the number of jumps within small time intervals and provide a theoretical error bound for this approximation. Additionally, we propose an Euler-Maruyama with restart solver, which achieves a provably lower error bound in estimating expected states and reduced variance compared to the standard solver. Experiments on both synthetic and real-world datasets demonstrate that Neural MJD consistently outperforms state-of-the-art deep learning and statistical learning methods. Our code is available at https://github.com/YuanpeiG/neural-MJD.
1UBC; 2Vector Institute; 3University of Texas at Austin; 4Canada CIFAR AI Chair
@@ 1 Introduction
Classical statistical models, e.g., Merton jump diffusion (MJD) [3] or more general Lévy processes [7], provide a principled approach for modeling such data with jumps. They are effective for small datasets with well-understood statistical properties [8, 9, 10]. However, their assumptions—such as independent and stationary increments—often fail in real-world non-stationary settings. Additionally, these models struggle to capture interdependencies across multiple time series, such as competition effects among colocated businesses [11] or spillover dynamics in stock markets driven by investor attention shifts [12, 13]. As a result, they are difficult to scale effectively to large datasets. In contrast, deep learning approaches have demonstrated strong empirical performance by effectively learning time-varying patterns from data [14, 15, 16, 17, 18, 19]. Despite their success, these models are often black-box in nature and lack explicit mathematical formulations to describe the underlying dynamics. This limits their interpretability and often results in poor generalization to non-stationary data with jumps.
@@ 3 Background
where $Q_{t}$ follows a compound Poisson process and captures the jump part. Specifically, $Q_{t}=\sum^{N_{t}}_{i=1}(Y_{i}-1)$, where $Y_{i}$ is the price ratio caused by the $i$-th jump event occurring at the time $t_{i}$, i.e., $Y_{i}={S_{t_{i}}}/{S_{t_{i}^{-}}}$ and $N_{t}$ is the total number of jumps up to time $t$. $S_{t_{i}}$ and $S_{t_{i}^{-}}$ are the prices after and before the jump at time $t_{i}$, respectively. $Y_{i}$ captures the relative price jump size since $\mathrm{d}S_{t_{i}}/S_{t_{i}}=(S_{t_{i}}-S_{t_{i}^{-}})/S_{t_{i}^{-}}=Y_{i}-1$. The price ratio $Y_{i}$ follows a log-normal distribution, i.e., $\ln Y_{i}\sim{\mathcal{N}}(\nu,\gamma^{2})$, where $\nu$ and $\gamma^{2}$ are the mean and the variance respectively. $N_{t}$ denotes the number of jumps that occur up to time $t$ and follows a Poisson process with intensity $\lambda$, which is the expected number of jumps per unit time. To make the expected relative price change $\mathbb{E}[\mathrm{d}S_{t_{i}}/S_{t_{i}}]$ remain the same as in the BS model in Eq. 1, MJD introduces an additional adjustment in the drift term of the diffusion, i.e., $-\lambda k\mathrm{d}t$ in Eq. 2. In particular, we have,
@@
where we use the assumption of MJD that “how much it jumps" (captured by $Y_{N_{t}}$) and “when it jumps" (captured by $N_{t}$) are independent. For the log-normal distributed $Y_{N_{t}}$, we can compute the expected jump magnitude $\mathbb{E}[Y_{N_{t}}-1]=\exp{(\nu+{\gamma^{2}}/{2})}-1$. To simplify the notation, we define $k\coloneqq\mathbb{E}[Y_{N_{t}}-1]$. For the Poisson process $N_{t}$, we have $\mathbb{E}[\mathrm{d}N_{t}]=\lambda\mathrm{d}t$. Therefore, we have $\mathbb{E}[\mathrm{d}Q_{t}]=\lambda k\mathrm{d}t$, which justifies the adjustment term $-\lambda k\mathrm{d}t$ in Eq. 2.
@@
The derivations of the above formulas are provided in App. A.
@@ 4 Methods
We consider the general time series prediction problem where, given the observed past data $\{S_{0},S_{-1},\dots,S_{-T_{p}}\}$, and optional contextual information (e.g., additional features) $C$, the goal is to predict the future values $\{S_{1},\dots,S_{T_{f}}\}$. Here $T_{p}$ and $T_{f}$ denote the past and future time horizons, respectively. While we use integer indices for denoting time, sampling time is not restricted to integers. Our model is built upon a diffusion process, which is inherently continuous in time and compatible with arbitrary sampling mechanism. An overview of our method is shown in Fig. 1.
@@ 4.1 Neural Non-Stationary Merton Jump Diffusion
Here $\lambda_{t}\in\mathbb{R}_{+}$ and $\sigma_{t},\mu_{t},k_{t}\in\mathbb{R}^{d}$, while $W_{t}$ denotes a $d$-dimensional standard Wiener process. With a slight abuse of notation, $\sigma_{t}\mathrm{d}W_{t}$ means element-wise product between two size-$d$ vectors. $N(\mathrm{d}t,\mathrm{d}y)$ is a Poisson random measure on $[0,T_{f}]\times\mathbb{R}^{d}$, which encodes both the timing and magnitude of jumps. Intuitively, $N(\mathrm{d}t,\mathrm{d}y)$ counts the number of jumps occurring within the infinitesimal time interval $[t,t+\mathrm{d}t]$ whose sizes fall within $[y,y+\mathrm{d}y]$. The jump component $S_{t}\int_{\mathbb{R}^{d}}(y-1)N(\mathrm{d}t,\mathrm{d}y)$ introduces abrupt discontinuities in the process, accounting for sudden shifts in data.
The statistical properties of the Poisson random measure are uniquely determined by its intensity measure $\lambda_{t}f_{Y}(t,y)\mathrm{d}y\mathrm{d}t$. In our model, the intensity measure controls time-inhomogeneous jump magnitudes and frequencies. Namely, jump times follow a Poisson process with time-dependent intensity $\lambda_{t}$ and jump magnitudes follow a time-dependent log-normal distribution, i.e., a jump $Y_{t}$ at time $t$ follows $\ln Y_{t}\sim\mathcal{N}(\nu_{t},\gamma_{t}^{2})$, where we denote the log-normal density of $Y_{t}$ as $f_{Y}(t,y)$. Similarly, we define $k_{t}\coloneqq\mathbb{E}[Y_{t}-1]=\exp(\nu_{t}+\frac{\gamma_{t}^{2}}{2})-1$ in the drift term. This makes Eq. 7 equivalent to using the compensated Poisson measure $\tilde{N}(\mathrm{d}t,\mathrm{d}y)\coloneqq N(\mathrm{d}t,\mathrm{d}y)-\lambda_{t}f_{Y}(t,y)\mathrm{d}y\mathrm{d}t$ to remove the expected contribution of jumps. Note Eq. 7 includes $k_{t},$ so that $\int_{\mathbb{R}^{d}}(y-1)\lambda_{t}f_{Y}(t,y)\mathrm{d}y\mathrm{d}t=\lambda_{t}\mathbb{E}[Y_{t}-1]\mathrm{d}t=\lambda_{t}k_{t}\mathrm{d}t.$ Namely, it can be rewritten as,
@@
where $f$ is a neural network parameterized by $\theta$. To simplify the notation, we denote the set of all observed data as ${\mathcal{C}}\coloneqq\{S_{0},S_{-1},\dots,S_{-T_{p}},C\}$ from now on.
The stochastic process described by the SDE in Eq. 7 is formally an instance of an additive process [61, Ch. 14], characterized by independent but non-stationary increments. If $\sigma_{t},\mu_{t}\in L^{2}$, i.e., they are square-integrable functions, and $\max_{\tau}(\int_{\mathbb{R}^{d}}|y^{2}|\lambda_{\tau}f_{Y}(\tau,y)\mathrm{d}y)<\infty$, then our non-stationary MJD has a unique solution for every $S_{0}>0$ [61, Theorem 14.1]. As our prediction time horizon is a closed domain $t\in[0,T_{f}]$, these conditions are easily satisfied as long as the neural network $f_{\theta}$ does not produce unbounded values. At any future time $T$, the explicit solution of the SDE is given by,
@@
Next, we model the conditional probability of log-return $\ln S_{t}$ given initial price $S_{0}$ and past data ${\mathcal{C}}$,
$P\left(\ln S_{T}|S_{0},{\mathcal{C}}\right)=\sum_{n=0}^{\infty}\frac{\exp(-\int_{0}^{T}\lambda_{t}\mathrm{d}t)}{n!}\Phi_{n},$ (10)
@@
Please refer to App. B for derivations. Evaluating Eq. 9 and Eq. 10 is non-trivial due to time inhomogeneity and jumps, typically requiring Monte Carlo methods or partial integro-differential equation techniques for approximate solutions [61, Ch. 6, Ch. 12].
@@ 4.2 Tractable Learning Method
While Eq. 10 provides the exact likelihood function, evaluating it precisely is impractical due to 1) integrals with time-dependent parameters lacking closed-form solutions and 2) the infinite series over the number of jumps. To learn the model via the maximum likelihood principle, we propose a computationally tractable learning objective with parameter bootstrapping and series truncation.
@@
[…]
where $a_{n,\delta}=\ln S_{t}+(\mu_{\rho_{t}}-\lambda_{\rho_{t}}k_{\rho_{t}}-{\sigma_{\rho_{t}}^{2}}/{2})\delta+n\nu_{\rho_{t}}$ and $b_{n,\delta}^{2}=\sigma^{2}_{\rho_{t}}\delta+\gamma^{2}_{\rho_{t}}n$. This approach eliminates the need for numerical simulation to compute the integrals in Eq. 10 and has been widely adopted for jump process modeling [62, 70].
@@
Derivation details are shown in App. B. Further, we jointly consider the likelihood of all future data:
@@
8:    Update $\theta$ via $-\nabla_{\theta}\sum_{\tau=1}^{T_{f}}\left(-\psi_{\tau}+\omega\|S_{\tau}-\hat{S}_{\tau}\|^{2}\right)$
@@
The training algorithm of our neural non-stationary MJD model is shown in Alg. 1. In computing the term $\ln P(\ln S_{\tau}|S_{\tau-1},{\mathcal{C}})$ of Eq. 13, instead of doing teacher forcing, we replace the ground truth $S_{\tau-1}$ with the conditional mean prediction $\mathbb{E}[S_{\tau-1}\mid{\mathcal{C}}]$ from Eq. 12. This design mitigates the discrepancy between training and inference while reducing error accumulation in sequential predictions, especially for non-stationary data. As shown in the ablation study in Sec. 5.3, this approach improves performance effectively. To further improve accuracy, we introduce an additional regularization term that encourages the conditional mean to remain close to the ground truth. Additionally, the for loop in Alg. 1 can be executed in parallel, as the conditional mean computation does not depend on sequential steps, significantly improving efficiency. Notably, our model imposes no restrictions on the neural network architecture, and the specific design details are provided in Sec. D.1.
Truncation error of likelihood function. Exact computation of $P(\ln S_{\tau}\mid S_{\tau-1},{\mathcal{C}})$ in Eq. 13 requires evaluating an infinite series, which is infeasible in practice. To address this, we truncate the series at order $\kappa\in\mathbb{N}_{+}$, i.e., limiting the maximum number of jumps between consecutive time steps. We establish the following theoretical result to characterize the decay rate of the truncation error:
@@
Let the likelihood approximation error in Sec. 4.2, truncated to at most $\kappa$ jumps, be
@@
The proof is provided in Sec. C.1. The truncation error is dominated by $\kappa$, with other time-dependent parameters absorbed into the big-$O$ notation. We set $\kappa$ to 5 to achieve better empirical performance.

### v2

@@ Abstract
While deep learning methods have achieved strong performance in time series prediction, their black-box nature and inability to explicitly model underlying stochastic processes often limit their robustness handling non-stationary data, especially in the presence of abrupt changes. In this work, we introduce Neural MJD, a neural network based non-stationary Merton jump diffusion (MJD) model. Our model explicitly formulates forecasting as a stochastic differential equation (SDE) simulation problem, combining a time-inhomogeneous Itô diffusion to capture non-stationary stochastic dynamics with a time-inhomogeneous compound Poisson process to model abrupt jumps. To enable tractable learning, we introduce a likelihood truncation mechanism that caps the number of jumps within small time intervals and provide a theoretical error bound for this approximation. Additionally, we propose an Euler-Maruyama with restart solver, which achieves a provably lower error bound in estimating expected states and reduced variance compared to the standard solver. Experiments on both synthetic and real-world datasets demonstrate that Neural MJD consistently outperforms state-of-the-art deep learning and statistical learning methods. Our code is available at https://github.com/DSL-Lab/neural-MJD.
1University of British Columbia; 2Vector Institute
3University of Texas at Austin; 4Canada CIFAR AI Chair
@@ 1 Introduction
Classical statistical models, e.g., Merton jump diffusion (MJD) [59] or more general Lévy processes [40], provide a principled approach for modeling such data with jumps. They are effective for small datasets with well-understood statistical properties [18, 7, 8]. However, their assumptions—such as independent and stationary increments—often fail in real-world non-stationary settings. Additionally, these models struggle to capture interdependencies across multiple time series, such as competition effects among colocated businesses [31] or spillover dynamics in stock markets driven by investor attention shifts [48, 29]. As a result, they are difficult to scale effectively to large datasets. In contrast, deep learning approaches have demonstrated strong empirical performance by effectively learning time-varying patterns from data [53, 74, 34, 47, 13, 58]. Despite their success, these models are often black-box in nature and lack explicit mathematical formulations to describe the underlying dynamics. This limits their interpretability and often results in poor generalization to non-stationary data with jumps. Notably, in previous deep learning time-series studies, “non-stationarity” typically refers to distributional shifts in the data over time. These works focus on mitigating such shifts using techniques like input-level normalization (e.g., DAIN [63], ST-norm [28], RevIN [42]), or domain adaptation (e.g., DDG-DA [50]). In contrast, our notion of “non-stationarity” centers on modeling a MJD process with parameters that evolve over time.
@@ 3 Background
where $Q_{t}$ follows a compound Poisson process and captures the jump part. Specifically, $Q_{t}=\sum^{N_{t}}_{i=1}(Y_{i}-1)$, where $Y_{i}$ is the price ratio caused by the $i$-th jump event occurring at the time $t_{i}$, i.e., $Y_{i}={S_{t_{i}}}/{S_{t_{i}^{-}}}$ and $N_{t}$ is the total number of jumps up to time $t$. $S_{t_{i}}$ and $S_{t_{i}^{-}}$ are the prices after and before the jump at time $t_{i}$, respectively. $Y_{i}$ captures the relative price jump size since $\mathrm{d}S_{t_{i}}/S_{t_{i}}=(S_{t_{i}}-S_{t_{i}^{-}})/S_{t_{i}^{-}}=Y_{i}-1$. The price ratio $Y_{i}$ follows a log-normal distribution, i.e., $\ln Y_{i}\sim{\mathcal{N}}(\nu,\gamma^{2})$, where $\nu$ and $\gamma^{2}$ are the mean and the variance respectively. $N_{t}$ denotes the number of jumps that occur up to time $t$ and follows a Poisson process with intensity $\lambda$, which is the expected number of jumps per unit time. To make the expected relative price change $\mathbb{E}[\mathrm{d}S_{t_{i}}/S_{t_{i}}]$ remain the same as in the BS model in Eq.˜1, MJD introduces an additional adjustment in the drift term of the diffusion, i.e., $\lambda k\mathrm{d}t$ in Eq.˜2. In particular, we have,
@@
where we use the assumption of MJD that “how much it jumps" (captured by $Y_{N_{t}}$) and “when it jumps" (captured by $N_{t}$) are independent. For the log-normal distributed $Y_{N_{t}}$, we can compute the expected jump magnitude $\mathbb{E}[Y_{N_{t}}-1]=\exp{(\nu+{\gamma^{2}}/{2})}-1$. To simplify the notation, we define $k\coloneqq\mathbb{E}[Y_{N_{t}}-1]$. For the Poisson process $N_{t}$, we have $\mathbb{E}[\mathrm{d}N_{t}]=\lambda\mathrm{d}t$. Therefore, we have $\mathbb{E}[\mathrm{d}Q_{t}]=\lambda k\mathrm{d}t$, which justifies the adjustment term $\lambda k\mathrm{d}t$ in Eq.˜2.
@@
The derivations of the above formulas are provided in App.˜A.
@@ 4 Methods
We consider the general time series prediction problem where, given the observed past data $\{S_{0},S_{-1},\dots,S_{-T_{p}}\}$, and optional contextual information (e.g., additional features) $C$, the goal is to predict the future values $\{S_{1},\dots,S_{T_{f}}\}$. Here $T_{p}$ and $T_{f}$ denote the past and future time horizons, respectively. While we use integer indices for denoting time, sampling time is not restricted to integers. Our model is built upon a diffusion process, which is inherently continuous in time and compatible with arbitrary sampling mechanism. An overview of our method is shown in Fig.˜1.
@@ 4.1 Neural Non-Stationary Merton Jump Diffusion
Here $\lambda_{t}\in\mathbb{R}_{+}$ and $\sigma_{t},\mu_{t},k_{t}\in\mathbb{R}^{d}$, while $W_{t}$ denotes a $d$-dimensional standard Wiener process. With a slight abuse of notation, $\sigma_{t}\mathrm{d}W_{t}$ means element-wise product between two size-$d$ vectors. $N(\mathrm{d}t,\mathrm{d}y)$ is a Poisson random measure on $0,T_{f}]\times\mathbb{R}^{d}$, which encodes both the timing and magnitude of jumps. Intuitively, a Poisson random measure extends the idea of a Poisson process to random events distributed over both time and magnitude spaces, and $N(\mathrm{d}t,\mathrm{d}y)$ counts the number of jumps occurring within the infinitesimal time interval $t,t+\mathrm{d}t]$ whose sizes fall within $y,y+\mathrm{d}y]$. The jump component $S_{t}\int_{\mathbb{R}^{d}}(y-1)N(\mathrm{d}t,\mathrm{d}y)$ introduces abrupt discontinuities in the process, accounting for sudden shifts in data.
The statistical properties of the Poisson random measure are uniquely determined by its intensity measure $\lambda_{t}f_{Y}(t,y)\mathrm{d}y\mathrm{d}t$. In our model, the intensity measure controls time-inhomogeneous jump magnitudes and frequencies. Namely, jump times follow a Poisson process with time-dependent intensity $\lambda_{t}$ and jump magnitudes follow a time-dependent log-normal distribution, i.e., a jump $Y_{t}$ at time $t$ follows $\ln Y_{t}\sim\mathcal{N}(\nu_{t},\gamma_{t}^{2})$, where we denote the log-normal density of $Y_{t}$ as $f_{Y}(t,y)$. Similarly, we define $k_{t}\coloneqq\mathbb{E}[Y_{t}-1]=\exp(\nu_{t}+\frac{\gamma_{t}^{2}}{2})-1$ in the drift term. This makes Eq.˜7 equivalent to using the compensated Poisson measure $\tilde{N}(\mathrm{d}t,\mathrm{d}y)\coloneqq N(\mathrm{d}t,\mathrm{d}y)-\lambda_{t}f_{Y}(t,y)\mathrm{d}y\mathrm{d}t$ to remove the expected contribution of jumps. Note Eq.˜7 includes $k_{t},$ so that $\int_{\mathbb{R}^{d}}(y-1)\lambda_{t}f_{Y}(t,y)\mathrm{d}y\mathrm{d}t=\lambda_{t}\mathbb{E}[Y_{t}-1]\mathrm{d}t=\lambda_{t}k_{t}\mathrm{d}t.$ Namely, it can be rewritten as,
@@
where $f$ is a neural network parameterized by $\theta$. To simplify the notation, we denote the set of all observed data as ${\mathcal{C}}\coloneqq\{S_{0},S_{-1},\dots,S_{-T_{p}},C\}$ from now on. Importantly, we only train a single network across all series and optimize the conditional log-likelihood as the training objective, which differs from standard statistical inference. At test time, the network produces context-dependent estimates $(\mu_{t},\sigma_{t},\lambda_{t},\nu_{t},\gamma_{t})$ across all future time in a single forward pass.
The stochastic process described by the SDE in Eq.˜7 is formally an instance of an additive process [79, Ch. 14], characterized by independent but non-stationary increments. If $\sigma_{t},\mu_{t}\in L^{2}$, i.e., they are square-integrable functions, and $\max_{\tau}(\int_{\mathbb{R}^{d}}|y^{2}|\lambda_{\tau}f_{Y}(\tau,y)\mathrm{d}y)<\infty$, then our non-stationary MJD has a unique solution for every $S_{0}>0$ [79, Theorem 14.1]. As our prediction time horizon is a closed domain $t\in[0,T_{f}]$, these conditions are easily satisfied as long as the neural network $f_{\theta}$ does not produce unbounded values. At any future time $T$, the explicit solution of the SDE is given by,
@@
Next, we model the conditional probability of log-return $\ln S_{t}$ given all the observed data $\mathcal{C}$},
$P\left(\ln S_{T}|{\mathcal{C}}\right)=\sum_{n=0}^{\infty}\frac{\exp(-\int_{0}^{T}\lambda_{t}\mathrm{d}t)}{n!}\Phi_{n},$ (10)
@@
Please refer to App.˜B for derivations. Evaluating Eq.˜9 and Eq.˜10 is non-trivial due to time inhomogeneity and jumps, typically requiring Monte Carlo methods or partial integro-differential equation techniques for approximate solutions [79, Ch. 6, Ch. 12].
@@ 4.2 Tractable Learning Method
While Eq.˜10 provides the exact likelihood function, evaluating it precisely is impractical due to 1) integrals with time-dependent parameters lacking closed-form solutions and 2) the infinite series over the number of jumps. To learn the model via the maximum likelihood principle, we propose a computationally tractable learning objective with parameter bootstrapping and series truncation.
@@
$\begin{aligned} P\left(\ln S_{t+\delta}|S_{t},{\mathcal{C}}\right)&=\sum_{n=0}^{\infty}P\left(\Delta N=n\right)P\left(\ln S_{t+\delta}|S_{t},{\mathcal{C}},\Delta N=n\right)\\ &=\sum_{n=0}^{\infty}\exp\left({-\lambda_{\rho_{t}}\delta}\right)\lambda_{\rho_{t}}^{n}\frac{\delta^{n}}{n!}\phi\left(\ln S_{t+\delta};a_{n,\delta},b^{2}_{n,\delta}\right),\end{aligned}$
where $a_{n,\delta}=\ln S_{t}+(\mu_{\rho_{t}}-\lambda_{\rho_{t}}k_{\rho_{t}}-{\sigma_{\rho_{t}}^{2}}/{2})\delta+n\nu_{\rho_{t}}$ and $b_{n,\delta}^{2}=\sigma^{2}_{\rho_{t}}\delta+\gamma^{2}_{\rho_{t}}n$. This approach eliminates the need for numerical simulation to compute the integrals in Eq.˜10 and has been widely adopted for jump process modeling [5, 26].
@@
Derivation details are shown in App.˜B. Further, we jointly consider the likelihood of all future data:
@@
8:  Update $\theta$ via $\nabla_{\theta}\sum_{\tau=1}^{T_{f}}\left(-\psi_{\tau}+\omega\|S_{\tau}-\hat{S}_{\tau}\|^{2}\right)$
@@
The training algorithm of our neural non-stationary MJD model is shown in Alg.˜1. In computing the term $\ln P(\ln S_{\tau}|S_{\tau-1},{\mathcal{C}})$ of Eq.˜13, instead of doing teacher forcing, we replace the ground truth $S_{\tau-1}$ with the conditional mean prediction $\mathbb{E}[S_{\tau-1}\mid{\mathcal{C}}]$ from Eq.˜12. This design mitigates the discrepancy between training and inference while reducing error accumulation in sequential predictions, especially for non-stationary data. As shown in the ablation study in Sec.˜5.3, this approach improves performance effectively. To further improve accuracy, we introduce an additional regularization term that encourages the conditional mean to remain close to the ground truth. Additionally, the for loop in Alg.˜1 can be executed in parallel, as the conditional mean computation does not depend on sequential steps, significantly improving efficiency. Notably, our model imposes no restrictions on the neural network architecture, and the specific design details are provided in Sec.˜D.1.
Truncation error of likelihood function. Exact computation of $P(\ln S_{\tau}\mid S_{\tau-1},{\mathcal{C}})$ in Eq.˜13 requires evaluating an infinite series, which is infeasible in practice. To address this, we truncate the series at order $\kappa\in\mathbb{N}_{+}$, i.e., limiting the maximum number of jumps between consecutive time steps. We establish the following theoretical result to characterize the decay rate of the truncation error:
@@
Let the likelihood approximation error in Sec.˜4.2, truncated to at most $\kappa$ jumps, be
@@
The proof is provided in Sec.˜C.1. The truncation error is dominated by $\kappa$, with other time-dependent parameters absorbed into the big-$O$ notation. We set $\kappa$ to 5 to achieve better empirical performance.